### v1

# State-witness contraction

Abstract
We present a method to construct entanglement witnesses for arbitrarily large multipartite systems, by tensoring and partial tracing existing states and witnesses. As a proof of principle we show that, using little shared quantum resources, the method allows to reuse witnesses unable to detect states with local positive partial transpositions into new ones able to do so. Moreover, we show that this technique allows to tailor both linear and nonlinear witnesses to specific states using semidefinite programming with comparatively low-dimensional variables. As an example, existing trace polynomial witnesses are significantly improved while preserving their symmetries and implementability with randomized measurements. Besides detecting entanglement, the method is shown to detect $k$-copy distillability. A recipe for the single-copy case is shown to be effective for generic and Werner states.
I Introduction
In the last decades, a large improvement has been done in the control over mutlipartite and high-dimensional quantum systems [1, 2]. This paves a road towards realistic quantum software [3], quantum networks [4, 5] and quantum devices where entanglement constitutes a major nonclassical resource. As a result, detecting the presence of entanglement in multipartite quantum systems becomes crucial [3, 6], but it is at the same time a theoretical and experimental challenge due to the growth of the state space with the local dimension [7, 8].
For this purpose, a vastly used technique is the evaluation of entanglement witnesses, namely observables whose negative expectation value detects the presence of entanglement [9, 10, 11]. However, important challenges remain: for some states, finding a witness detecting them requires unaccessible computational resources [12, 13]; finding new entanglement witnesses for multipartite systems of interest is challenging [14, 15, 16]; and adjusting a witness to a state cannot be done systematically in general [7]. A promising tool detecting larger sets of states [17] are nonlinear witnesses acting on multiple copies [18, 19]. While these overcome some limitations of linear witnesses [17, 20], problems like tailoring nonlinear witnesses to states are even more challenging than in the linear case.
Here we provide a new method to derive new witnesses using existing states and witnesses (Prop. 1 and Fig. 1). We show that existing witnesses can be lifted to larger ones with further detection capabilities (Observation 2). Naturally, the construction given in this work allows to partially adjust specific witnesses to desired states efficiently – see Eq. (4). As a very special case, the method allows to construct multicopy witnesses, which are shown to detect $k$-copy distillability under some conditions. We compare this method with trace polynomial multicopy witnesses we previously introduced in [20], over which significant advantage is shown with little computational effort.
II State-witness contraction
The following result provides a systematic recipe to construct new entanglement witnesses from existing states and witnesses (depicted in Fig. 1). A deeper analysis in terms of positive maps in the special case of trace polynomials is given in [21].
Proposition 1 (State-witness contraction).
Let $\{R^{(j)}\}_{j=1}^{n}$ be $k_{j}$-partite states or witnesses. Let $\tau$ be a $n$-partite state or witness. Then, the operator
[…]
has a nonnegative expectation value, $\operatorname{tr}(\mathcal{W}_{\tau}\varrho)\geq 0$, on separable $(k_{1}+...+k_{n}-n)$-partite states $\varrho$.
This can be seen from the duality between the block-positive cone and the cone of separable states [22] as follows.
Proof.
Using the coordinate-free definition of the partial trace, for any state $\sigma^{(j)}_{k}$ and product state $\varrho^{(j)}=\varrho_{1}^{(j)}\otimes...\otimes\varrho_{k-1}^{(j)}$ we have $0\leq\operatorname{tr}(R^{(j)}\cdot\varrho^{(j)}\otimes\sigma^{(j)}_{k})=\operatorname{tr}(\theta^{(j)}\otimes\sigma^{(j)}_{k})$, where we define $\theta^{(j)}:=\operatorname{tr}_{1...k-1}(R_{j}\cdot\varrho_{1}^{(j)}\otimes...\otimes\varrho_{k-1}^{(j)}\otimes\mathds{1})$. Equivalently, $\theta^{(j)}$ is positive semidefinite for all $1\leq j\leq n$. Therefore, for any $n$-partite state or witness $\tau$, we have the inequality $0\leq\operatorname{tr}(\theta^{(1)}\otimes...\otimes\theta^{(n)}\cdot\tau)=\operatorname{tr}(\mathcal{W}_{\tau}\varrho)$. Here the equality is obtained by plugging in the definition (1) of $\mathcal{W}_{\tau}$, and we assume $\varrho:=\varrho^{(1)}\otimes...\otimes\varrho^{(n)}$ where each $\varrho^{(j)}$ is a product $(k_{j}-1)$-partite state. By convexity, $\operatorname{tr}(\mathcal{W}_{\tau}\varrho)\geq 0$ also holds for any separable $(k_{1}+...+k_{n}-n)$-partite state $\varrho$.
Figure 1: State-witness contraction yielding a four-partite witness from two three-partite operators. Here we take the tensor product of a witness $W$ and a positive operator $M$. Then, instead of obtaining a multicopy or trace polynomial witness as in [20, 23], the first (or last, equivalently) subsystems $A_{1}$ and $B_{1}$ are contracted (multiplied and partial traced) with a state or witness $\tau_{A_{1}B_{1}}$ to be chosen – see Eqs. (1) and (3). This allows to lift $W$ to a larger witness $\mathcal{W}_{\tau}$ with further detection capabilities (Observation 2), which moreover can be partially optimized over $\tau$ to detect specific four-partite states – see Eq. (4).
III Locally bound entanglement
As a case of interest, we will employ this method to reuse existing highly symmetric witnesses given in [23], which belong to a larger class of methods involving permutations in the Hilbert space [24, 25]. While these are desired due to their symmetries and implementation, their use in detecting states with local positive partial transpositions is a current challenge [26, 14, 23, 27, 28, 29, 30, 31]. In particular, consider the witness
$W=\frac{P_{2,1}}{4}-P_{1^{3}}\,$ (2)
constructed from the so-called Young projectors $P_{2,1}=\eta_{d}\big{(}2(\operatorname{id})-(123)-(132)\big{)}/6$ and $P_{1^{3}}=\eta_{d}\big{(}(\operatorname{id})-(12)-(13)-(23)+(123)+(132)\big{)}/6$, where $\eta_{d}(\pi)\mathinner{|v_{1}\rangle}\mathinner{|v_{2}\rangle}\mathinner{|v_{3}\rangle}=\mathinner{|v_{\pi^{-1}(1)}\rangle}\mathinner{|v_{\pi^{-1}(2)}\rangle}\mathinner{|v_{\pi^{-1}(3)}\rangle}\in{\mathds{C}^{d}}^{\otimes 3}$ represents the permutation $\pi\in S_{3}$ onto ${(\mathds{C}^{d})^{\otimes 3}}$. This witness can be derived from immanant inequalities [32, 33] using the techniques in [23, 26, 20]. Linear optimization shows that it can detect three-partite states with zero, one or two positive partial transpositions, but not three (see Table II in [21]). Using Proposition 1 with an ancillary Bell state $\mathinner{|\phi^{+}\rangle}=\sum_{i=0}^{2}\mathinner{|ii\rangle}/\sqrt{3}$, it can be recycled to detect four-partite states whose local parties have positive partial transpositions, and thus cannot be distilled with local operations with classical communication if the four parties are spatially separated. This is a type of unlockable bound entanglement [34, 35, 36, 37, 38, 39], which we refer to as locally bound.
Observation 2.
Let $W$ be the witness of Eq. (2), and let $P:=P_{1^{3}}$ be the three-qutrit antisymmetrizer. Choose $\tau=|\phi^{+}\rangle\langle\phi^{+}|$. The four-qutrit entanglement witness
$\mathcal{W}_{\tau}=\operatorname{tr}_{A_{3}B_{3}}\Big{(}W_{A_{1}A_{2}A_{3}}\otimes P_{B_{1}B_{2}B_{3}}\cdot\mathds{1}\otimes\tau_{A_{3}B_{3}}\Big{)}$ (3)
can detect four-partite entangled states $\varrho_{A_{1}A_{2}B_{1}B_{2}}$ with $\varrho^{T_{i}}\geq 0$ for $i\in\{A_{1},A_{2},B_{1},B_{2}\}$. In particular, $\min\{\operatorname{tr}(\mathcal{W}_{\tau}):\varrho^{T_{i}}\geq 0\}\approx-0.000276$
Note that this can be achieved using the shared quantum resources provided by the Bell state in an ancillary system $A_{3}B_{3}$. Yet, these can be reduced significantly: using a less entangled state $\mathinner{|\phi_{s}\rangle}=\sqrt{s}\mathinner{|00\rangle}+\sqrt{1-s}\mathinner{|11\rangle}$ embedded into $\mathds{C}^{3}\otimes\mathds{C}^{3}$ with $s=0.9999$, a minimum value $\min_{\varrho^{T_{i}}\geq 0}\langle\mathcal{W}_{\tau}\rangle_{\varrho}\approx-6.051\cdot 10^{-5}$ is obtained using only $-s\log_{2}(s)-(1-s)\log_{2}(1-s)\approx 0.0015$ ebits of entanglement.
IV Witnesses tailored to states
To some degree, witnesses constructed in Proposition 1 can be efficiently adjusted to specific states: consider a fixed set of $n$ $k_{j}$-partite witnesses (or states) $R^{(i)}$ and a target $(k_{1}+...+k_{n}-n)$-partite state $\varrho$ in local dimension $d$ to be detected. Let us choose $\tau$ to be optimized over all possible states and decomposable witnesses [9],
$\tau^{*}=\text{argmin}_{\tau}\quad$ $\operatorname{tr}\bigg{(}\bigotimes_{i=1}^{n}R^{(i)}\cdot\varrho\otimes\tau\bigg{)}$ (4)
so that $\tau=X+\sum_{\mathcal{S}\subset\{k_{1},...,k_{n}\}}X_{\mathcal{S}}$
$X,X_{\mathcal{S}}^{T_{\mathcal{S}}}\geq 0$
$\operatorname{tr}(\tau)=1.$
Here $\mathcal{S}$ denotes any possible subset of the last subsystems $\{k_{1},...,k_{n}\}$. While by Proposition 1 one can choose $\tau$ to be any block-positive operator (any state or witness), the optimization becomes accessible with semidefinite programming if $\tau$ is restricted to be a decomposable witness or a state [40]. Notice that the variable $\tau$ is of size $d^{n}$ [41], even though the target state to be detected is described by a density matrix of size $d^{(k-1)n}$ (assuming $k_{1}=...=k_{n}:=k$ for simplicity).
As a simple example, using that $\operatorname{tr}(\phi^{+}\cdot X\otimes Y)=\operatorname{tr}(X^{T}Y)$ and $\operatorname{tr}\big{(}\eta_{d}(12)\cdot X\otimes Y\big{)}=\operatorname{tr}(XY)$ where $\phi^{+}=|\phi^{+}\rangle\langle\phi^{+}|$ and $\eta_{d}(12)$ is the SWAP operator, a witness given by
$\mathcal{W}_{\tau^{*}}=\operatorname{tr}_{A_{2}B_{2}}\Big{(}\eta_{d}(12)_{A_{1}B_{1}}\otimes\phi^{+}_{B_{1}B_{2}}\cdot\varrho_{A_{1}B_{1}}\otimes\tau^{*}_{A_{2}B_{2}}\Big{)}$ (5)
with $\tau^{*}=\text{argmin}_{\tau}\big{\{}\operatorname{tr}(\mathcal{W}_{\tau}\varrho):\tau\geq 0,\operatorname{tr}\tau=1\big{\}}$ detects any state $\varrho$ with nonpositive partial transpose. Similarly, standard witness tailoring techniques using diagonalization [9] are recovered with this example: the duality between positive maps and witnesses [42] is used to detect any state detected by a positive map $\Lambda_{A}\otimes\operatorname{id}_{B}(\varrho_{AB})\not\geq 0$, replacing $\phi^{+}$ in Eq. (5) for $\Lambda^{\dagger}\otimes\operatorname{id}(\eta_{d}(12))$ where $\Lambda^{\dagger}$ satisfies $\operatorname{tr}(\Lambda(X)Y)=\operatorname{tr}(X\Lambda^{\dagger}(Y))$ for all operators $X$ and $Y$.
V Multiple copies
Here we will focus on the case where the state to be detected consists of multiple copies $\varrho^{\otimes k}$ of a target state $\varrho$. We shall demonstrate that the technique introduced in this work allows to significantly improve our previous trace-polynomial multicopy constructions given in [20]. The latter are obtained as a very special case of the constructions given by Proposition 1, by setting the input state to be composed of $k-1$ copies of a $n$-partite target state $\varrho^{\otimes k-1}$. For example, consider a Bell state affected by white noise,
$\varrho=p|\phi^{+}\rangle\langle\phi^{+}|+(1-p)\frac{\mathds{1}}{d^{2}}\,,$ (6)
which is often used as a worse-case scenario in noisy setups [43]. In [20] we showed that this state can be detected for a range of values of $p$ with the nonlinear witness $\mathcal{W}=(\mathds{1}-6P_{1^{3}})_{A_{123}}\otimes(P_{1^{3}})_{B_{123}}$ acting on three copies of $\varrho$, through $0<\operatorname{tr}(\mathcal{W}\varrho^{\otimes 3})=\operatorname{tr}(\mathcal{W}_{\varrho}\varrho^{\otimes 2})$ where we denote $\mathcal{W}_{\varrho}:=\operatorname{tr}_{A_{3}B_{3}}(\mathcal{W}\cdot\mathds{1}\otimes\varrho_{A_{3}B_{3}})$ for easier comparison to our present work. From Proposition 1 it becomes clear that $\mathcal{W}$ can be better tuned to $\varrho$ through Eq. (4), in this case leading to (Fig. 1)
$\mathcal{W}_{\tau}:=\operatorname{tr}_{A_{3}B_{3}}\Big{(}\mathcal{W}\cdot(\mathds{1}_{A_{1}A_{2}B_{1}B_{2}}\otimes\phi^{+}_{A_{3}B_{3}})\Big{)}\,.$ (7)
In Table 1 we compare the smallest values of $p$ so that states of Eq. (6) are detected with $\mathcal{W}_{\varrho}$ (obtained in [20]) and the optimized witness $\mathcal{W}_{\phi^{+}}$ in (7).
Figure 2: Tailoring multicopy witnesses to states. Linear witnesses detect regions of entangled states separated by hyperplanes (red area) through $\langle W\rangle_{\varrho}=\operatorname{tr}(W\varrho)<0$. Multicopy witnesses $\mathcal{W}$ detect regions of entangled states through $\langle\mathcal{W}\rangle_{\varrho^{\otimes k}}<0$ [18, 19], separated by curved hypersurfaces (magenta and blue areas). In particular, $\mathcal{W}$ can be constructed by tensoring linear witnesses [20]. By replacing one of the copies by a variable state or witness $\tau$, the state-witness contraction technique (SWC) allows to systematically detect a larger set of states (green area –see Table 1 for an example) and approach further the set of separable states (SEP).
Note that this method of tailoring witnesses of the form proposed in [20] preserves their invariant properties:
Observation 3.
Let $\mathcal{W}_{\tau}$ in Eq. (4) be constructed from trace polynomial operators $R^{(j)}$ obtained by immanant inequalities according to [23]. Then, the expression $\operatorname{tr}(\mathcal{W}_{\tau}\varrho^{\otimes k-1})$ is invariant under local unitaries acting on $\varrho$.
Proof.
Local unitaries act as $U^{\otimes k_{j}-1}$ onto the first $k_{j}-1$ copies of each $j$’th local party. A minimum obtained for $\varrho_{AB}$ by $\tau_{AB}$ is also obtained for $U\otimes V\varrho_{AB}U^{\dagger}\otimes V^{\dagger}$ by $U\otimes V\tau_{AB}U^{\dagger}\otimes V^{\dagger}$ (extension to the multipartite case is straightforward). Thus local unitaries act in practice as $U^{\otimes k_{j}}$ on the $j$’th party. The rest of the argument is the same as in [20], namely $U^{\otimes k_{j}}$ commutes with the representation $\eta_{d}(\pi)$ of permutations $\pi\in S_{k_{j}}$ due to the Schur-Weyl duality. ∎
$p^{*}$:  detect $d=3$ $d=4$ $d=5$ $d=6$
$\langle\mathcal{W}\rangle_{\varrho^{\otimes^{3}}}<0$ [20] $0.727$ $0.674$ $0.630$ $0.594$
$\langle\mathcal{W}_{\phi^{+}}\rangle_{\varrho^{\otimes^{2}}}<0$ (4) $0.319$ $0.262$ $0.222$ $0.193$
$\varrho\in\text{ENT}$    $0.25$ $0.2$ $0.1\overline{6}$ $0.\overline{142857}$
Table 1: Advantage of tailored nonlinear witnesses, over non-tailored. In each column we display the threshold parameter $p^{*}$ so that the sate (6) is detected for all values $p\geq p^{*}$, obtained numerically, with different methods for local dimension $d=3,4,5,6$. We compare these values corresponding to the non-tailored witness $\mathcal{W}$ we provided in [20] (first row, $\tau=\varrho$); and the tailored version of this witness (7) (second row, $\tau=|\phi^{+}\rangle\langle\phi^{+}|$). In this toy model, the state is entangled (ENT) for $p>(d+1)^{-1}$ and separable otherwise [43].
Besides being unitary invariant, the witnesses we constructed in [20] can be implemented with randomized measurements [24, 44, 25, 45]. They key idea is that since only moments of the density matrix $\varrho$ are involved in the expectation values, then postprocessing the classical shadows can be done by matrix multiplication, thus avoiding large arrays. Note that the improvement in Eq. (4) preserves this implementability: one replaces each $k$’th classical shadow $\varrho^{r_{k}}=\bigotimes_{i=1}^{n}\varrho^{k}_{i}$ obtained from a set of random outcomes $r_{k}$ by the known optimal state $\tau$, and operates analogously to approximate $\operatorname{tr}(\mathcal{W}_{\tau}\varrho^{\otimes k-1})$.
VI Distillability
By the Choi-Jamiołkowski isomorphism, each state or witness opertor $R_{j}$ can be seen as the Choi operator of a map $\Psi_{j}(\varrho_{j}^{T}):=\operatorname{tr}_{1...k_{j}-1}(R^{(j)}\cdot{\varrho^{j}}\otimes\mathds{1}_{k})$ from the space of operators acting on $\mathds{C}^{d_{1}^{j}}\otimes...\otimes\mathds{C}^{d_{k_{j}-1}}$ to the space of operators acting on $\mathds{C}^{d_{k_{j}}}$, whose output is positive semidefinite if the inputs are products of positive semidefinite operators (further details are given in [21]). In the bipartite case, consider $R^{(1)}=P_{A}\Pi^{\otimes k}$ and $Q_{B}:=R^{(2)}=Q_{B}\Pi^{\otimes k}$ where $P_{A}$ and $Q_{B}$ are projectors onto a aubspace of $\mathds{C}^{d_{1}^{j}}\otimes...\otimes\mathds{C}^{d_{k_{j}}}$, and $\Pi$ is a local projector onto $\mathds{C}^{2}$ (see Tab. I and Fig. 2 in [21] for specific results derived in this case). Then the action of $\Psi_{A_{1}...A_{k-1}}\otimes\Psi_{B_{1}...B_{k-1}}(\varrho_{AB}^{\otimes k-1})$ projects $k-1$ copies to the two-qubit subspace $\mathds{C}^{2}\otimes\mathds{C}^{2}$, where where the PPT-criterion detects both entanglement and distillability [46, 43].
The framework introduced in this work naturally provides a criterion for $k$-copy distillability, since a state is distillable if and only if it can be transformed into a two-qubit entangled state through local projections on $k$ copies [47, 48]. This shall be illustrated in specific examples involving a single copy, considering the two-qudit antisymmetrizer $P_{1,1}=(\eta_{d}()-\eta_{d}(12))/2$ and a two-dimensional local projector $\Pi$ diagonal in the computational basis: we demonstrate that most generic states in a random sample are detected to be one-copy distillable by using $R^{(1)}=P_{1,1}\Pi^{\otimes 2}$ (see [21] for a detailed numerical analysis), effectively detecting the qubit-qudit distillable entanglement present in a subspace of a two-qudit system. Moreover, choosing $R^{(1)}=R^{(2)}=P_{1,1}\Pi^{\otimes 2}$, the operator $\mathcal{W}=R^{(1)}_{A_{1}A_{2}}\otimes R^{(2)}_{B_{1}B_{2}}$ detects both entanglement and one-copy distillability on all Werner states $\varrho_{W}$ known to be distillable, namely
$\varrho_{W}=pP_{2,0}^{(d)}/d_{s}+(1-p)P_{1,1}^{(d)}/d_{a}$ (8)
within $p\in(p_{0},1/2]$ with $p_{0}=\frac{d+1}{4d-2}$ [49], where $d_{s}=d(d+1)/2$ and $d_{s}=d(d-1)/2$. Indeed, one verifies that
$\operatorname{tr}_{A_{1}B_{1}}\Big{(}\mathcal{W}\cdot{\varrho_{W}}_{A_{1}B_{1}}\otimes\mathds{1}_{A_{2}B_{2}}\Big{)}=pP_{a}^{(2)}/d_{s}+(1-p)P_{s}^{(2)}/d_{a}\,.$ (9)
Therefore, choosing the two-qubit SWAP operator $\tau=2|\phi^{+}\rangle\langle\phi^{+}|^{T_{A}}=\eta_{2}(12)$, for $p\in(p_{0},1/2]$ we have $\operatorname{tr}(\mathcal{W}_{\tau}\varrho_{W})<0$.
VII Conclusions
We have developed a novel technique to design new entanglement witnesses, and demonstrated its capabilities with examples. As a proof of concept we have shown that the method allows to use witnesses unable to detect locally bound entangled states, to construct new larger witnesses able to do so. Moreover, the witnesses constructed in this way can be naturally be optimized to desired states to be detected. This can be done with semidefinite programming whose variable size is smaller than the state to be detected. As a very special case, the method applies to construct nonlinear witnesses acting on multiple copies. With little computational resources, we have seen that our previous multicopy construction provided in [20] can be tuned to significantly better detect noisy entangled states, while the desired properties of unitary invariance and implementability with randomized measurements have been shown to be preserved. Moreover, the technique introduced leads to practical recipes to detect distillability, which are demonstrated to be effective for relevant families of states.
The examples provided demonstrate that this technique opens an avenue to constructing both linear and nonlinear witnesses with desired properties, systematically tailored for a variety of purposes. In particular, the efficient implementability of our method relies on optimizing over a variable $\tau$ restricted to be a state or a witness that is decomposable across some bipartition. For further research, it would be interesting to analyze whether fully non-decomposable witnesses can be obtained when the operator $\tau$ is chosen to be fully non-decomposable; and whether optimal witnesses in the sense of [50] can be systematically obtained.
Acknowledgements:   The author is thankful to Felix Huber, Robin Krebs, Ferran Riera-Sàbat, Some Sankar Bhattacharya and Anna Sanpera for comments and support. Support by the Foundation for Polish Science through TEAM-NET project POIR.04.04.00-00-17C1/18-00 and by NCN QuantERA Project No. 2021/03/Y/ST2/00193 is gratefully acknowledged. The author also acknowledges financial support from Spanish MICIN (projects: PID2022:141283NBI00;139099NBI00) with the support of FEDER funds, the Spanish Goverment with funding from European Union NextGenerationEU (PRTR-C17.I1), the Generalitat de Catalunya, the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project -Quantum Spain Project- through the Recovery, Transformation and Resilience Plan NextGeneration EU within the framework of the Digital Spain 2026 Agenda.